The combined age of a mother and her twin daughters is 58 years. The mother was 25 year old when the twins were born. Write and solve an equation to find the age of each of the three people.

Found 2 solutions by stanbon, DrBeeee:
Answer by stanbon(75887)

The combined age of a mother and her twin daughters is 58 years. The mother was 25 year old when the twins were born. Write and solve an equation to find the age of each of the three people.
-----------
Equation:
m + 2d = 58
m = d + 25
----
Substitute for "m" and solve for "d":
d+25 + 2d = 58
3d = 33
d = 11 (Each of the twin daughters is 11)
---
Solve for "m":
m = d + 25
m = 11 + 25 = 36 ( mom's age now)
================
Cheers,
Stan H.
================

Answer by DrBeeee(684)

Let m = mom's age
Let t = twin's age
Sum of ages = 58
Since mom was 25 when twins were born
m = t + 25
sum is
t + t + m = 58
2t + m = 58
2t + t + 25 = 58
3t + 25 = 58
3t = 58 - 25
3t = 33
t = 33/3
t = 11
m = t +25
m = 36
Is (t+t+m = 58)?
Is (11+11+36=58)?
Is (58=58)? Yes
Answer: The twins are 11 and mother is 36.